### v2

# Design of a CsI(Tl) Calorimeter for Muonium-to-Antimuonium Conversion Experiment

Abstract
The Muonium-to-Antimuonium Conversion Experiment (MACE) is proposed to search for charged lepton flavor violation and increase the sensitivity by three orders of magnitude compared to the PSI experiment in the 1990s. A clear signature of this conversion is the positron produced from antimuonium decay. This paper presents a near-$4\pi$-coverage calorimeter designed for MACE, which can provide an energy resolution of 9% at 511 keV, and a signal efficiency of 55.5% for annihilation $\gamma$-ray events at maximum. Detailed Monte-Carlo simulations using MACE offline software based on Geant4 are performed for geometry optimization, coincidence system design, background estimation, and benchmark detector validation.
Muonium-to-antimuonium conversion, Charged lepton flavor violation, Electromagnetic calorimeter, Inorganic scintillator detector
I Introduction
In the Standard Model (SM), lepton flavor violation is strictly forbidden. However, the discovery of neutrino oscillations has shown that neutral leptons can violate this conservation law, hinting at the existence of new physics beyond the Standard Model [1, 2]. Consequently, the Charged Lepton flavor Violation (CLFV) is naturally expected within the framework of new physics. Muon, the lightest massive unstable charged lepton with a reasonably long lifetime, serves as an ideal probe in precision frontier experiments aimed at searching for CLFV processes. Since the 1940s, several CLFV experiments with muon have been conducted, consistently reporting increasingly stringent upper limits over the decades. Key channels of interest include $\mu^{+}\rightarrow e^{+}\gamma$, $\mu^{+}\rightarrow e^{+}e^{-}e^{+}$, and $\mu^{-}N\rightarrow e^{-}N$ [3]. In recent years, notable experiments searching for these CLFV channels include MEG and its upgrade, MEG-II [4], Mu3e [5], COMET [6], and Mu2e [7]. However, the processes mentioned above violate lepton flavor by only one unit.
Muonium (M) is an exotic atom consisting of a bound state of $\mu^{+}e^{-}$, while its antiatom, antimuonium ($\bar{\text{M}}$), is composed of $\mu^{-}e^{+}$. From the perspective of the Standard Model Effective Field Theory (SMEFT), the spontaneous M-to-$\bar{\text{M}}$ conversion ($\mu^{+}e^{-}\rightarrow\mu^{-}e^{+}$), which violates lepton flavor by two units instead of one, can probe distinct effective operators and thereby constrain a different parameter space [8, 9, 10]. As a result, the key innovation in investigating M-to-$\bar{\text{M}}$ conversion is providing additional constraints on specific new physics models, distinct from other CLFV processes [11, 12, 13]. The most recent measured upper limit ($8.3\times 10^{-11}$ at 90% C.L.) for the conversion probability in a 0.1 T magnetic field was reported by the MACS experiment at PSI in 1998 [14, 15]. Another experiment focused on M-to-$\bar{\text{M}}$ conversion was proposed and conceptually designed at PRISM in 2003, as reported in [16], but did not progress to further studies. There is also a proposal at J-PARC in Japan to search for this process using a novel scheme [17]. Surprisingly, the limit of M-to-$\bar{\text{M}}$ conversion has remained unchallenged for two decades. There is an urgent need for a new-generation experiment.
In light of significant advancements in muon beamlines and particle detector techniques nowadays, the Muonium-to-Antimuonium Conversion Experiment (MACE) is proposed to search for M-to-$\bar{\text{M}}$ conversion and has the potential to improve the upper limit by three orders of magnitude [18]. With the construction of the China advanced NUclear physics research Facility (CNUF) [19], MACE is expected to be the first muon-CLFV experiment in China, providing extra perspectives compared to other concerned experiments. The proposed MACE detector consists of a Michel electron Magnetic Spectrometer (MMS), a Microchannel Plate (MCP) [20, 21], and an Electromagnetic Calorimeter (ECAL) to probe the conversion signal through final state coincidence. A schematic of the MACE apparatus is shown in Fig. 1.
Figure 1: Schematic view of the MACE detector with a simulated M-to-$\bar{\text{M}}$ signal.
In the MACE concept, surface muons with a momentum of 26.3 MeV$/c$ are stopped at a production target where M is formed and then possibly diffuses into the vacuum and converts into $\bar{\text{M}}$. The decay products of $\bar{\text{M}}$ include an electron with Michel energy (up to 52.8 MeV) and a positron with atomic-shell energy ($\sim 13.5$ eV). The Michel electron will be reconstructed by MMS which is expected to achieve $10^{-4}$ level suppression of background in the ECAL end through preliminary simulation. On the other hand, the atomic-shell positron will be accelerated by an electric field and guided through the solenoid with transverse-position conservation to MCP, where the positron annihilates into two back-to-back 511 keV $\gamma$-rays. The cylindrical magnet with a field strength of 0.1 T outside ECAL is used to constrain the positron beam spot after transportation. As a rare-process searching experiment, MACE requires high sensitivity for the annihilation $\gamma$-rays to discriminate the signal from backgrounds and retain as much event information as possible. Thus, an ECAL with good energy resolution and high efficiency is essential for the design of MACE.
[…]
The main purpose of this study is to propose a design for the MACE $4\pi$ calorimeter using the Monte-Carlo simulation method. In this work, the geometry of ECAL is generated and optimized using the MACE offline software based on the Geant4 toolkit [35, 36]. Subsequently, the energy resolution, signal efficiency, and background level of ECAL are estimated, providing a reference for the MACE conceptual design. This paper is structured as follows Section I introduces the MACE calorimeter design and geometry optimization. Section II presents the simulation results for both signal and background. Section III describes the experiment setups of a benchmark detector to validate the energy resolution study in the MACE calorimeter design. The paper concludes with a discussion of the conclusions and perspectives for future work in Section IV.
II Design and Simulation of the MACE Calorimeter
II.1 Overview
To meet the requirements of enhanced sensitivity (e.g., excellent energy resolution, high detection efficiency, and good geometrical acceptance), we propose a near-$4\pi$-coverage polyhedral calorimeter design. In this section, we present a design scheme for ECAL, including details on scintillation materials and photosensors. This proposed design will serve as the basis for the simulation work discussed in the following sections.
The $4\pi$ ECAL composed of polyhedral crystals is geometrically referred to as the Goldberg polyhedron in mathematics, which was first introduced by Michel Goldberg in 1937 [37]. The Goldberg polyhedron consists of pentagons and hexagons with every three faces meeting at the same vertex. It is similar in appearance to the sphere as an inpolyhedron and possesses an icosahedral rotational symmetry. We generate the Goldberg polyhedron mesh using the following steps:
  1. (1)
    Generate a regular icosahedron mesh inscribed on the unit sphere;
  2. (2)
    Apply the Loop subdivision algorithm [38] to the mesh and project new vertices to the sphere;
  3. (3)
    Project the centroids of each face onto the sphere and connect them to form a Goldberg polyhedron mesh.
  4. (4)
    Scale the mesh radially to the desired size.
The Polygon Mesh Processing Library (PMP) [39] is used to implement the above algorithm, which generates the ECAL geometry in the MACE offline software. A Class I GP$(8,0)$ Goldberg polyhedron mesh inscribed in the unit sphere is produced, consisting of a total of 642 faces, including 12 regular pentagons and 630 irregular hexagons, which can be categorized into 9 types. By scaling the obtained mesh using the ECAL inner radius $\boldsymbol{r}$ and the crystal length $\boldsymbol{l}$, a parameterized ECAL geometry which can be easily adjusted is established in the MACE offline software via G4TessellatedSolid in Geant4. Modules created from the different types of faces are identified as Type-PEN and Type-HEX (01–09). Fig. 2 displays the ECAL geometry rendered using Unreal engine [40].
(a) 3D rendered image of the crystals.
(b) An event display for ECAL.
Figure 2: Schematic view of the ECAL geometry.
Calorimeters with inorganic scintillation crystals generally provide excellent energy resolution and thus have been widely used in particle physics experiments for decades. Commonly used scintillators in calorimeters include NaI(Tl), pure CsI, CsI(Tl), BGO, PbWO4, and others [25]. Table 1 summarizes the properties of these inorganic scintillators. NaI(Tl) was widely used in the past decades due to its low cost and high light output. Nonetheless, it is hygroscopic and has a relatively low density, making it less suitable for large-scale particle physics experiments at present. Despite its short radiation length resulting in higher energy deposition efficiency, the low light output of BGO and PbWO4 leads to insufficient energy resolution. Although the pure CsI possesses a relatively fast decay time, its low light output leads to poor energy resolution. In the MACS experiment, the energy resolution of the pure CsI calorimeter for single $\gamma$ event and double $\gamma$ events was 69% and 52%, respectively [41]. Doping thallium (Tl) in pure CsI can significantly increase its light yield and result in excellent energy resolution. The MACE ECAL is required to maintain high precision for $e^{+}e^{-}$ annihilation $\gamma$-rays. Consequently, CsI(Tl) is chosen as the reference scintillator in this ECAL design.
Table 1: Properties of several inorganic scintillation crystals [42].
[…]
  • $f$
    fast component
  • $s$
    slow component
The photosensor is also a crucial component in the calorimeter system. Its technical specifications (e.g., sensitive area, photon detection efficiency) determine the efficiency with which optical photons are converted into electrical signals that can ultimately be recorded. Commonly used photosensors in current particle and nuclear physics experiments include Photomultiplier Tubes (PMTs) and Silicon Photomultipliers (SiPMs) [43]. PMTs are relatively less expensive compared to SiPMs for the same sensitive area. Despite their limited quantum efficiency, high operating voltage, and shielding requirements in magnetic fields, PMTs are widely used in experiments. On the other hand, SiPMs offer excellent photon detection efficiency, a compact size, and tolerance to magnetic fields. However, they come with higher costs, dark current due to semiconductor thermal noise, and crosstalk between adjacent pixels [44]. Currently, PMTs are considered the benchmark photosensor in this paper.
II.2 Optimization of Geometry Parameters
The ECAL geometry is generated by expanding the Goldberg polyhedron mesh radially, with the ECAL inner radius and crystal length determining the module size. In a scintillator-based detection system, the energy resolution is primarily determined by the scintillation light yield and its collection efficiency. Additionally, the geometric dimensions of each module are crucial for light collection efficiency [45], considering factors such as the $\gamma$-ray energy deposition rate and the propagation of optical photons within the crystal, which are ultimately collected by the photosensor. To enhance the sensitivity of ECAL, the energy resolution should be optimized by scanning geometry parameters, including the crystal length and the ECAL inner radius.
A detailed Monte-Carlo simulation of the module Type-HEX06, which is characterized by both the largest number and size, is performed using MACE offline software based on Geant4 v11.2.2. The physics processes are modeled with the reference physics list FTFP_BERT_EMZ, G4OpticalPhysics, and the Unified surface model. Additionally, the optical properties of Teflon-wrapped CsI(Tl) crystals (such as light yield, emission spectrum, decay time, etc.) and the PDE of the photosensor are considered based on manufacturer data sheets. Monoenergetic 511 keV $\gamma$-rays from a point-biased source are directed onto the center of a crystal optically coupled with a photosensor, which is registered as a Sensitive Detector (SD). The area of SD is adjusted to fully cover the outer surface of a Type-HEX06 crystal to estimate the intrinsic energy resolution of the crystal. The crystal length varies from 10 cm to 19 cm, and the inner radius of ECAL varies from 15 cm to 24 cm. Peak finding is performed using TSpectrum within the ROOT framework [46] to record the maximum count of scintillation photons spectra collected by the SD.
Meanwhile, the detection efficiency is given by the ratio of the number of events with energy greater than 500 keV to the total number of incident particles. The results of scanning each geometry parameter combination are detailed in Fig. 3.
(a) Results of photon counts.
(b) Results of detection efficiency.
(c) Results of F.o.M.
Figure 3: Scanning results of collected photon counts at maximum and detection efficiency of Type-HEX06 module with different geometry parameters including crystal length and ECAL inner radius. A deeper color suggests a better performance in all figures.
It should be noted that there is a trade-off between energy resolution and detection efficiency. The result demonstrates that increasing of the crystal volume causes a decrease in light collection efficiency and a deterioration in energy resolution, as investigated by Refs. [47, 48]. It is reasonable to assume that, as optical photons travel a longer distance, some of them will be absorbed by the material directly or after multiple reflections, resulting in reduced light collection efficiency. On the other hand, the crystal length has a limited effect on the detection efficiency because the longitudinal leakage of 511 keV $\gamma$-rays is negligible. Incident particles are more likely to escape laterally due to Compton scattering; thus, detection efficiency increases with the ECAL inner radius, which affects the module width. To balance the energy resolution and the detection efficiency, a Figure-of-Merit (F.o.M) function is constructed as follows:
[…]
where $N_{pho}$ is the maximum count of scintillation photons spectra, $\eta$ is the detection efficiency. The parameters are ultimately chosen to be a crystal length of 10 cm and an ECAL inner radius of 20 cm based on consideration of both F.o.M and spatial constraints.
In the context of the MACE experiment, the signal positrons from $\bar{\text{M}}$ decay are transported through a solenoid into the ECAL chamber. Some modules must be removed to ensure that the incoming signal positrons are not obstructed. Specifically, the beam windows are set with an upstream radius of 50 mm and a downstream radius of 5 mm. Consequently, 19 modules at the beam entrance and 1 module at the exit are removed, resulting in a total of 622 modules and achieving 97% coverage of the $4\pi$ solid angle.
Figure 4: Calibration by Geant4 simulation of Type-HEX06 module. The red line is the FWHM curve fit by simulation data (blue dots), while the green line is the fit energy resolution.
The class G4RadioactiveDecayPhysics is used to simulate the calibration of the Type-HEX06 module with radioactive sources, including 22Na, 137Cs, and 152Eu. The crystal is coupled with a 1.5” diameter PMT. An isotropic point source, wrapped in 1 mm thick plastic, is placed at the entrance of the module. The energy dependence of FWHM is fit using  Eq. 2:
$\text{FWHM}=a+b\sqrt{E_{0}+cE_{0}^{2}}~{},$ (2)
where the best-fit value of $a$, $b$, $c$ are -1.56, 1.82, $1.72\times 10^{-4}$, respectively. The fit FWHM and energy resolution results of each full-energy peak are plotted in Fig. 4. For Type-HEX06 modules, the energy resolution is 9% at 511 keV, which is noticeably better than the resolution achieved by the calorimeter used in the MACS experiment [41]. Future improvements could include the adoption of large-area SiPM arrays with higher PDE.
II.3 Signal and Backgrounds
II.3.1 M-to-$\bar{\text{M}}$ Conversion Signal
In the M-to-$\bar{\text{M}}$ conversion process, $\bar{\text{M}}$ decays into a Michel electron and an atomic-shell positron. The positron is then accelerated electrostatically and transported through the solenoid into the ECAL chamber, where it is detected by the MCP. Due to the low kinetic energy of the positron, it will annihilate in MCP, producing a pair of back-to-back $\gamma$-rays. The expected signal for the positron in M-to-$\bar{\text{M}}$ conversion event will be identified through a coincidence of MCP and ECAL responses. This section details the coincidence studies conducted using the MACE offline software, which provides insights into event reconstruction and sensitivity evaluation.
A beam of $10^{8}$ $\mu^{+}$ events with a momentum of 26.3 MeV$/c$ entering the MACE detector has been simulated. The $\mu^{+}$ forms M in the production target with a certain probability, which then completely converts into $\bar{\text{M}}$ which subsequently decays. Event sampling is performed according to the detection efficiency curve of MCP, and positron events with multiple hits on MCP are filtered out. After a positron annihilates, the produced $\gamma$-rays either interact with the sensitive volume of ECAL or escape the chamber. The coincidence between MCP and ECAL responses ensures that the annihilation signals originate from the same M-to-$\bar{\text{M}}$ conversion event. Following the coincidence, a clustering algorithm is applied to search seed modules and their adjacent modules. The energy and direction information of the hits is reconstructed, and only two simultaneous responses that satisfy certain criteria will be selected. The event selection criteria can be summarized as follows:
  • Threshold cut: $E_{thr}>0.2~{}\text{MeV}$;
  • Energy difference cut: $\Delta E_{recon}<0.1~{}\text{MeV}$;
  • Angle cut: $\theta_{recon}>160^{\circ}$.
These criteria stand for trigger threshold, energy difference, and angle of two reconstructed hits, respectively. Finally, the ECAL response spectra are smeared using the FWHM curve described by Eq. 2. Events that fall within the $3\sigma$ interval of the full-energy peak are identified as signal events.
Figure 5: Spectra of original and reconstructed annihilation signals of M-to-$\bar{\text{M}}$ conversion by simulation.
Fig. 5 summarizes the simulated spectra of the positron annihilation signal. In the original $\gamma$ spectrum, the 511 keV full-energy peak accounts for only about 30% of the total counts. This is attributed to some of the $\gamma$-rays scattering to adjacent modules after depositing part of their energy, which forms a Compton continuum. Additionally, a peak at approximately 80 keV appears in the single $\gamma$ spectrum. This peak is attributed to characteristic X-rays emitted by lead atoms in the MCP. As a result, the signal efficiency is significantly reduced. However, by clustering the energy responses from each event, the total energy of the scattered photons can be reconstructed. As can be seen from Fig. 5, the energy has been reconstructed effectively. The expected signal is thus represented by the $3\sigma$ interval of 1.022 MeV.
Table 2: Contributions of signal Efficiency.
Items Efficiency
Geometry 97%
No cut 55.5%
Energy difference cut 55.3%
Threshold cut 49.9%
Angle cut 43.5%
Total 42.8%
The result of the simulation indicates a 42.8%-55.5% signal efficiency in $3\sigma$ for ECAL. Losses of signal efficiency are as detailed in Table 2. Applications of various event selection criteria will depend on the real background level. The energy resolution and signal efficiency of ECAL are both superior to those of the MACS calorimeter, which is expected to significantly enhance the sensitivity of MACE.
II.3.2 Beam Background
The muon beam, produced by stopping $\pi^{+}$ on the surface of a target, often includes positrons or other particles with similar momentum [49]. These background particles can enter the ECAL chamber in the same manner as the signal positrons, potentially causing accidental coincidences between MCP and ECAL. Therefore, it is crucial to estimate their potential contribution.
A parallel simulation involving $10^{11}$ $e^{+}$ events was conducted for comparison with the signal. No coincidence events were observed in this simulation. According to the technical specifications of the Experimental Muon Source (EMuS) proposed at CSNS [50], the optimal ratio of positrons to muons is expected to be 0.1% [51]. The beam positron background level is estimated based on this data. Assuming a $\mu^{+}$ beam flux of $10^{8}/$s, and a coincidence probability of $10^{-4}$ with the magnetic spectrometer, the upper limit of the background rates is established at 0.007/0.007/0.007 /year at 90% C.L. Thus, the ECAL is expected to be tolerant of both continuous and pulsed beams.
II.3.3 Cosmic-Ray Muon Background
Cosmic-ray muons are another source of background in ECAL. The EcoMug generator [52] is integrated into the MACE offline software to simulate cosmic-ray muon events on various types of surfaces (plane, cylinder, and half-sphere) with accurate angle and momentum distributions.
To evaluate cosmic-ray muon background level in a potential future environment for the MACE detector, a hypothetical 18 m deep tunnel has been modeled in Geant4. A concrete volume, represented by G4Box with dimensions $50\times 50\times 18~{}\text{m}^{3}$, is created, with an air-filled cylindrical region, represented by G4Tubs, subtracted from its bottom to form a tunnel with a 1 m radius and 10 m length. ECAL is positioned at the center of the tunnel, surrounded by a copper solenoid and an iron shield. Cosmic-ray muons are emitted from a $50\times 50~{}\text{m}^{2}$ plane source at sea level. With a rate of 129 Hz, cosmic-ray muon events equivalent to 1 hour of exposure are generated.
The simulation results indicate that cosmic-ray muons still contribute to the background in the ECAL even after passing through the concrete and shielding. This includes some events where $e^{+}e^{-}$ annihilation produces 511 keV $\gamma$ rays. After applying coincidence selection, no events remain, resulting in background rates of less than 2/2/2 /year when considering coincidences with the Magnetic Spectrometer. Typically, a veto system can be installed outside ECAL to further reject cosmic-ray muon background [6, 53].
III Validation with a Benchmark Detector
III.1 Experiment Setups and Results
To characterize the parameters of the CsI(Tl) scintillator, a benchmark detector was constructed and calibrated using various radioactive sources at Sun Yat-sen University. It consists of a $3\times 3\times 8~{}\text{cm}^{3}$ CsI(Tl) crystal and a 1.5” PMT (HAMAMATSU CR284 with E2183-500 socket assembly). The crystal is wrapped with ESR (glued) and coupled with PMT using optical grease, finally coated with 3M tape (Fig. 6b). The readout system uses a Lecroy HDO4054A oscilloscope (Fig. 6c) for signal triggering, integration, and digitization.
(a) CsI(Tl) crystal.
(b) Benchmark detector.
(c) Lecroy HDO4054A with a typical waveform of CsI(Tl) on screen.
(d) CAEN V1730SB digitizer and the VMEbus.
Figure 6: Photographs of the experiment setups.
Figure 7: Measured spectra of ${}^{137}\text{Cs}$ (in green), ${}^{152}\text{Eu}$ (in red) at SYSU and ${}^{22}\text{Na}$ (in blue) at IMP.
The benchmark detector is calibrated using ${}^{137}\text{Cs}$ and ${}^{152}\text{Eu}$ radioactive sources. Radioactive sources were positioned close to the detector, and data acquisition was performed for 1800 s. When the PMT pulse signal exceeds the threshold, the oscilloscope triggers, integrates the waveform and records the charge value. The calibration used the 662 keV peak from ${}^{137}\text{Cs}$ and the 40 keV, 122 keV, 244 keV, 344 keV, and 1408 keV peaks from ${}^{152}\text{Eu}$, as shown in Fig. 7. The spectra display the full-energy peaks of both isotopes. Using the FWHM curve, the energy resolution at 511 keV is calculated to be 11.5%.
A more precise experiment was conducted to investigate the energy resolution at the Institute of Modern Physics, Chinese Academy of Sciences in Lanzhou, China. The absolute gain of the CR284 PMT was accurately calibrated in advance using single photoelectron (SPE) response measurements. A waveform generator sends pulses with a voltage of 2.65 V, a frequency of 10 kHz, and a width of 5 ns to an LED to produce weak light emission. A 120 s measurement scans the high voltage ranging from 850 V to 1250 V. The PMT response waveforms are read out and recorded in ROOT files using the CAEN V1730SB digitizer (Fig. 6d), which is externally triggered by the waveform generator. After offline pulse processing, the SPE peaks are fit to determine their mean charge value which is then converted to the absolute gain.
Table 3: Energy resolution (FWHM) performance at 511 keV of the benchmark detector with various reflectors.
Reflector Energy resolution
Teflon 10.3%
ESR (non-glued) 11.1%
ESR (glued) 11.5%
Tyvek 13.1%
Subsequently, the energy resolution performance with various reflectors is tested. The CsI(Tl) crystal is wrapped with materials including Teflon, ESR (non-glued), and Tyvek. Other experiment setups were the same as Fig. 6b, with ${}^{22}\text{Na}$ and ${}^{137}\text{Cs}$ sources used. Measurements were conducted for 600 s with the PMT operating at 850 V. The energy resolution is directly calculated from the FWHM, and the results are listed in Table 3.
III.2 Comparison and Discussion
It can be seen from Table 3 that Teflon provides a better energy resolution compared to other materials. This indicates that the properties of reflector material have a significant impact on the energy resolution. Ref. [54] measured the reflectivity of commonly used reflectors. For the tested reflectors in Table 3, the reflectivity values are 0.99, 0.985, and 0.97, respectively. It can be inferred that with the Teflon reflector, the PMT collects more photons due to its higher reflectivity leading to improved energy resolution. The energy resolution with glued ESR in the earlier experiment was worse compared to the resolution with non-glued ESR obtained later. This is due to the absorption or refraction of light as it passes through the glue. Moreover, the air gap between the non-glued reflectors and the crystal enhances the collection efficiency of scintillation photons through total internal reflection. In summary, Teflon is a promising reflector material for enhancing the energy resolution of MACE ECAL.
This study primarily relies on Geant4 simulation for the design of the MACE calorimeter, requiring verification of the reliability of various physical processes in the simulation and a comparison with experimental results. To test the reliability of the simulation results discussed in Section II, we conducted a detailed simulation of the benchmark CsI(Tl) detector using Geant4 with the same physics processes. The CsI(Tl) crystal, Teflon reflector, optical grease, and CR284 PMT volume are modeled identically to the experimental setup. At the end of the simulation, the number of produced photoelectrons is recorded in a ROOT file event by event. The energy resolution is 8.2% at 511 keV in the simulation, while the experiment result is 10.3% (see Table 3).
Figure 8: The photoelectron spectra of ${}^{137}\text{Cs}$ in experiment and simulation.
The difference in energy resolution between the experiment and the simulation is reasonable. Fig. 8 shows the spectra of photoelectrons reconstructed from experimental data and those recorded in simulation using ${}^{137}\text{Cs}$ source. Considering the mean photoelectron number of the 662 keV peak, the Geant4 simulation results are generally consistent with the experiment. However, there are still minor, predictable discrepancies. Several possible explanations are as follows. Firstly, the PMT gain value measurement may contain deviations caused by electronic noise or statistical errors. As a result, the number of photoelectrons reconstructed from the waveform is biased by approximately 76 PEs compared with the simulation. Secondly, the ideal simulation did not account for the natural radioactive background, resulting in a lack of response in the higher energy region. However, this contribution is expected to be limited due to the coincidence and reconstruction.
The comparison between the experiment and the simulation of the benchmark detector gives the expected results with little divergence, which validates the design of the MACE ECAL and provides a foundation for further research.
IV Conclusions and Perspectives
After more than two decades of development of high-intensity muon source and detector technique, the first proposed muon-CLFV experiment in China, MACE, is expected to enhance the sensitivity of M-to-$\bar{\text{M}}$ conversion by three orders of magnitude. This study focuses on the offline-software-based design of the $4\pi$ ECAL geometry for MACE. The geometry parameters were optimized using the Monte-Carlo simulation method, and the performance such as energy resolution, signal efficiency, and background level were evaluated. Simulation results indicate an energy resolution of 9% at 511 keV, and a signal efficiency of 55.5% for annihilation $\gamma$-ray events at maximum, significantly outperforming the MACS calorimeter design. This research establishes a foundation for the subsequent technical development of the ECAL and supports the conceptual design of MACE.
However, the studies on accidental coincidence backgrounds from the beam and cosmic rays are currently idealized, and the effects of electronic noise and natural radioactive backgrounds have not yet been evaluated. A prototype of ECAL will be constructed for additional validation in progress in these areas. Further studies are anticipated to evaluate the performance enhancement of ECAL with large-area SiPM array readout. Upgrades of ECAL are expected with the integration of an inner tracker system. Moreover, future works will also focus on developing systems for cosmic-ray vetoing, electronic readout, trigger logic, and reconstruction algorithms.
V Acknowledgements
The authors would like to acknowledge the MACE Working Group for their invaluable teamwork. We also extend our gratitude to the colleagues at the SMOOTH lab, as well as to Ke Gong and Chenfeng Yang at Sun Yat-sen University, for their fruitful discussions and assistance related to this work. The simulation benefited greatly from the provision of computing resources by the National Supercomputer Center in Guangzhou. We are also grateful to the Southern Center for Nuclear-Science Theory (SCNT) at the Institute of Modern Physics in the Chinese Academy of Sciences for hospitality.

### v3

@@
 Abstract
-The Muonium-to-Antimuonium Conversion Experiment (MACE) is proposed to search for charged lepton flavor violation and increase the sensitivity by three orders of magnitude compared to the PSI experiment in the 1990s.
+The Muonium-to-Antimuonium Conversion Experiment (MACE) is proposed to search for charged lepton flavor violation and increase the sensitivity by more than two orders of magnitude compared to the MACS experiment at PSI in 1999.
 A clear signature of this conversion is the positron produced from antimuonium decay.
-This paper presents a near-$4\pi$-coverage calorimeter designed for MACE, which can provide an energy resolution of 9% at 511 keV, and a signal efficiency of 55.5% for annihilation $\gamma$-ray events at maximum.
+This paper presents a near-$4\pi$-coverage calorimeter designed for MACE, which can provide an energy resolution of 10.8% at 511 keV, and a signal efficiency of 78.3% for annihilation $\gamma$-ray events.
 Detailed Monte-Carlo simulations using MACE offline software based on Geant4 are performed for geometry optimization, coincidence system design, background estimation, and benchmark detector validation.
@@
 As a result, the key innovation in investigating M-to-$\bar{\text{M}}$ conversion is providing additional constraints on specific new physics models, distinct from other CLFV processes [11, 12, 13].
-The most recent measured upper limit ($8.3\times 10^{-11}$ at 90% C.L.) for the conversion probability in a 0.1 T magnetic field was reported by the MACS experiment at PSI in 1998 [14, 15].
-Another experiment focused on M-to-$\bar{\text{M}}$ conversion was proposed and conceptually designed at PRISM in 2003, as reported in [16], but did not progress to further studies.
+The most recent measured upper limit ($8.3\times 10^{-11}$ at 90% C.L.) for the conversion probability in a 0.1 T magnetic field was reported by the MACS experiment at PSI in 1999 [14, 15].
+Another experiment focused on M-to-$\bar{\text{M}}$ conversion was proposed and conceptually designed at PRISM in 2003, as reported in [16].
 There is also a proposal at J-PARC in Japan to search for this process using a novel scheme [17].
+But none of these have progressed to further studies.
 Surprisingly, the limit of M-to-$\bar{\text{M}}$ conversion has remained unchallenged for two decades.
 There is an urgent need for a new-generation experiment.
-In light of significant advancements in muon beamlines and particle detector techniques nowadays, the Muonium-to-Antimuonium Conversion Experiment (MACE) is proposed to search for M-to-$\bar{\text{M}}$ conversion and has the potential to improve the upper limit by three orders of magnitude [18].
-With the construction of the China advanced NUclear physics research Facility (CNUF) [19], MACE is expected to be the first muon-CLFV experiment in China, providing extra perspectives compared to other concerned experiments.
-The proposed MACE detector consists of a Michel electron Magnetic Spectrometer (MMS), a Microchannel Plate (MCP) [20, 21], and an Electromagnetic Calorimeter (ECAL) to probe the conversion signal through final state coincidence.
+In light of significant advancements in muon beamlines and particle detector techniques nowadays, the Muonium-to-Antimuonium Conversion Experiment (MACE) is proposed to search for M-to-$\bar{\text{M}}$ conversion and has the potential to improve the upper limit by three orders of magnitude [18, 19].
+With the construction of the China advanced NUclear physics research Facility (CNUF) and other local muon sources [20, 21, 22, 23], MACE is expected to be the first muon-CLFV experiment in China, offering unique perspectives compared to other related experiments, including those at colliders.
+The proposed MACE detector consists of a Michel electron Magnetic Spectrometer (MMS), a Positron Transport System (PTS), a Microchannel Plate (MCP) [24, 25], and an Electromagnetic Calorimeter (ECAL) to probe the conversion signal through final state coincidence.
+The specifications of muon beamline in MACE is expected to produce a surface muon flux on target of $10^{8}~{}\mu^{+}/$s, with a repetition rate of 30–50 kHz and a pulse width duty cycle below 20%.
 A schematic of the MACE apparatus is shown in Fig.
@@
 In the MACE concept, surface muons with a momentum of 26.3 MeV$/c$ are stopped at a production target where M is formed and then possibly diffuses into the vacuum and converts into $\bar{\text{M}}$.
-The decay products of $\bar{\text{M}}$ include an electron with Michel energy (up to 52.8 MeV) and a positron with atomic-shell energy ($\sim 13.5$ eV).
+The decay products of $\bar{\text{M}}$ include a Michel electron with an energy up to 52.8 MeV and a positron with atomic-shell energy ($\langle E_{\text{k}}\rangle=13.5$ eV).
 The Michel electron will be reconstructed by MMS which is expected to achieve $10^{-4}$ level suppression of background in the ECAL end through preliminary simulation.
-On the other hand, the atomic-shell positron will be accelerated by an electric field and guided through the solenoid with transverse-position conservation to MCP, where the positron annihilates into two back-to-back 511 keV $\gamma$-rays.
+On the other hand, the atomic-shell positron will be accelerated by an electric field and guided through the PTS.
+The transverse position of the $e^{+}$ is conserved during the transportation to MCP, with a flight time FWHM of 6.9 ns.
 The cylindrical magnet with a field strength of 0.1 T outside ECAL is used to constrain the positron beam spot after transportation.
+While the MCP, with a timing resolution of 500 ps, is hit by a positron, the ECAL will be triggered and measure the produced two back-to-back annihilation $\gamma$-rays.
 As a rare-process searching experiment, MACE requires high sensitivity for the annihilation $\gamma$-rays to discriminate the signal from backgrounds and retain as much event information as possible.
@@
 Subsequently, the energy resolution, signal efficiency, and background level of ECAL are estimated, providing a reference for the MACE conceptual design.
-This paper is structured as follows Section I introduces the MACE calorimeter design and geometry optimization.
+This paper is structured as follows: Section I introduces the MACE calorimeter design and geometry optimization.
 Section II presents the simulation results for both signal and background.
-Section III describes the experiment setups of a benchmark detector to validate the energy resolution study in the MACE calorimeter design.
+Section III describes the experiment setups of a benchmark detector, specifically the assembly used for an initial test of the CsI(Tl) crystal.
 The paper concludes with a discussion of the conclusions and perspectives for future work in Section IV.
@@
 II.1 Overview
-To meet the requirements of enhanced sensitivity (e.g., excellent energy resolution, high detection efficiency, and good geometrical acceptance), we propose a near-$4\pi$-coverage polyhedral calorimeter design.
-In this section, we present a design scheme for ECAL, including details on scintillation materials and photosensors.
-This proposed design will serve as the basis for the simulation work discussed in the following sections.
+To meet the requirements of enhanced sensitivity (e.g., an excellent energy resolution, a high detection efficiency, and a good geometric acceptance), we propose a near-$4\pi$-coverage polyhedral calorimeter design.
+In this section, we present a design scheme for ECAL, including details on the selection of scintillation materials and photosensors.
+This proposed design will serve as a basis for the simulation work discussed in the following sections.
 The $4\pi$ ECAL composed of polyhedral crystals is geometrically referred to as the Goldberg polyhedron in mathematics, which was first introduced by Michel Goldberg in 1937 [37].
@@
     Scale the mesh radially to the desired size.
+(a) 3D rendered image of the crystals.
+(b) An event display for ECAL.
+Figure 2: Schematic view of the ECAL geometry.
 The Polygon Mesh Processing Library (PMP) [39] is used to implement the above algorithm, which generates the ECAL geometry in the MACE offline software.
@@
 2 displays the ECAL geometry rendered using Unreal Engine [40].
-(a) 3D rendered image of the crystals.
-(b) An event display for ECAL.
-Figure 2: Schematic view of the ECAL geometry.
+Table 1: Properties of several inorganic scintillation crystals [41, 42]
+[…]
+  • f fast component, s slow component
 Calorimeters with inorganic scintillation crystals generally provide excellent energy resolution and thus have been widely used in particle physics experiments for decades.
@@
 Table 1 summarizes the properties of these inorganic scintillators.
-NaI(Tl) was widely used in the past decades due to its low cost and high light output.
+The selection of scintillation crystal for MACE ECAL is based on the energy region of interest less than 1 MeV.
+NaI(Tl) was widely used in the past decades due to its low cost and high scintillation light output.
 Nonetheless, it is hygroscopic and has a relatively low density, making it less suitable for large-scale particle physics experiments at present.
-Despite its short radiation length resulting in higher energy deposition efficiency, the low light output of BGO and PbWO4 leads to insufficient energy resolution.
-Although the pure CsI possesses a relatively fast decay time, its low light output leads to poor energy resolution.
-In the MACS experiment, the energy resolution of the pure CsI calorimeter for single $\gamma$ event and double $\gamma$ events was 69% and 52%, respectively [41].
-Doping thallium (Tl) in pure CsI can significantly increase its light yield and result in excellent energy resolution.
-The MACE ECAL is required to maintain high precision for $e^{+}e^{-}$ annihilation $\gamma$-rays.
+The problem of hygroscopy is identical for LaBr3(Ce).
+Moreover, considering the overall budget of this experiment, ultrafast and bright crystals such as LYSO(Ce), GAGG(Ce), or LaBr3(Ce) are not currently included but may be considered for potential future upgrades.
+The low light output of PbWO4 leads to insufficient energy resolution in the low-energy regime.
+The same is true for pure CsI.
+In the MACS experiment, the energy resolution of the pure CsI calorimeter for single $\gamma$ event and double $\gamma$ events was 69% and 52%, respectively [44, 14].
+Doping thallium (Tl) in pure CsI can significantly increase its light yield, making it one of the brightest scintillators known.
+The objective of MACE ECAL is is to achieve excellent energy resolution for $\gamma$-rays produced in $e^{+}e^{-}$ annihilation.
+This requires optimizing the light yield of the scintillator, which is a key factor in precise energy measurements.
 Consequently, CsI(Tl) is chosen as the reference scintillator in this ECAL design.
-Table 1: Properties of several inorganic scintillation crystals [42].
-[…]
-  • $f$
-    fast component
-  • $s$
-    slow component
 The photosensor is also a crucial component in the calorimeter system.
@@
 Commonly used photosensors in current particle and nuclear physics experiments include Photomultiplier Tubes (PMTs) and Silicon Photomultipliers (SiPMs) [45].
-PMTs are relatively less expensive compared to SiPMs for the same sensitive area.
-Despite their limited quantum efficiency, high operating voltage, and shielding requirements in magnetic fields, PMTs are widely used in experiments.
-On the other hand, SiPMs offer excellent photon detection efficiency, a compact size, and tolerance to magnetic fields.
+PMTs are relatively more cost-effective than SiPMs for the same sensitive area but are constrained by the low quantum efficiency (QE), high operating voltage, and shielding requirements in magnetic fields.
+On the other hand, SiPMs offer excellent photon detection efficiency (PDE), a compact size, and tolerance to magnetic fields.
 However, they come with higher costs, dark current due to semiconductor thermal noise, and crosstalk between adjacent pixels [46].
 Currently, PMTs are considered the benchmark photosensor in this paper.
+A preliminary test of the SiPM performance is also discussed in Section III.2.
 II.2 Optimization of Geometry Parameters
@@
 A detailed Monte-Carlo simulation of the module Type-HEX06, which is characterized by both the largest number and size, is performed using MACE offline software based on Geant4 v11.2.2.
-The physics processes are modeled with the reference physics list FTFP_BERT_EMZ, G4OpticalPhysics, and the Unified surface model.
+The module mentioned above and below is defined as a combination of a scintillation crystal and a photosensor assembled together.
+The physics processes are modeled with the reference physics list QBBC_EMZ, G4OpticalPhysics, and the Unified surface model.
 Additionally, the optical properties of Teflon-wrapped CsI(Tl) crystals (such as light yield, emission spectrum, decay time, etc.) and the PDE of the photosensor are considered based on manufacturer data sheets.
+Detailed configurations of the simulation are presented in Table 2.
+Note that to better simulate the energy resolution of CsI(Tl), a Fano Factor $F_{\text{N}}$ is introduced [48, 49], and its squared value is assigned to the RESOLUTIONSCALE parameter in Geant4.
 Monoenergetic 511 keV $\gamma$-rays from a point-biased source are directed onto the center of a crystal optically coupled with a photosensor, which is registered as a Sensitive Detector (SD).
@@
 3.
+Table 2: Specification of material properties.
+Material Density Refractive index Absorption length Reflectivity
+$\rho$ (g$/\text{cm}^{3}$) $n$ $\lambda$ (cm) $R$
+Crystal CsI(Tl) 4.51 1.79 37 -
+Reflector Teflon 2.25 1.35 - 0.98
+Coupler EJ-550 0.97 1.46 100 -
+PMT window Glass 2.40 1.49 -
+SiPM window Epoxy 1.18 1.55
+PMT cathode Bialkali 2.00 - 0
+SiPM sensor Silicon 2.33
 (a) Results of photon counts.
@@
 Incident particles are more likely to escape laterally due to Compton scattering; thus, detection efficiency increases with the ECAL inner radius, which affects the module width.
-To balance the energy resolution and the detection efficiency, a Figure-of-Merit (F.o.M) function is constructed as follows:
+To balance the energy resolution and the detection efficiency, a Figure-of-Merit (F.o.M) function is constructed, following the function used in Ref.
+[53]:
 […]
-where $N_{pho}$ is the maximum count of scintillation photons spectra, $\eta$ is the detection efficiency.
-The parameters are ultimately chosen to be a crystal length of 10 cm and an ECAL inner radius of 20 cm based on consideration of both F.o.M and spatial constraints.
-In the context of the MACE experiment, the signal positrons from $\bar{\text{M}}$ decay are transported through a solenoid into the ECAL chamber.
+where $N_{\text{pho}}$ is the maximum count of scintillation photons spectra, $\eta$ is the detection efficiency.
+These two quantities are normalized by their mean values.
+The geometric parameters are ultimately chosen to be a crystal length of 10 cm and an ECAL inner radius of 20 cm based on consideration of both F.o.M and spatial constraints.
+In the context of the MACE experiment, the signal positrons from $\bar{\text{M}}$ decay are transported through a solenoid into the region of positron detection system.
 Some modules must be removed to ensure that the incoming signal positrons are not obstructed.
@@
 Consequently, 19 modules at the beam entrance and 1 module at the exit are removed, resulting in a total of 622 modules and achieving 97% coverage of the $4\pi$ solid angle.
-Figure 4: Calibration by Geant4 simulation of Type-HEX06 module. The red line is the FWHM curve fit by simulation data (blue dots), while the green line is the fit energy resolution.
-The class G4RadioactiveDecayPhysics is used to simulate the calibration of the Type-HEX06 module with radioactive sources, including 22Na, 137Cs, and 152Eu.
-The crystal is coupled with a 1.5” diameter PMT.
+For an estimation of energy resolution, the class G4RadioactiveDecayPhysics is used to simulate the calibration of the Type-HEX06 module with radioactive sources, including 22Na, 137Cs, and 152Eu.
+Instead of the fully-covered SD in pervious simulation, the crystal is coupled with an 1.5-inch diameter PMT volume.
 An isotropic point source, wrapped in 1 mm thick plastic, is placed at the entrance of the module.
@@
 2:
-$\text{FWHM}=a+b\sqrt{E_{0}+cE_{0}^{2}}~{},$ (2)
-where the best-fit value of $a$, $b$, $c$ are -1.56, 1.82, $1.72\times 10^{-4}$, respectively.
+$\text{FWHM}=a+b\sqrt{E_{0}}~{},$ (2)
+where the best-fit value of $a$, $b$ are $-7.4707$ and $2.7638$, respectively.
 The fit FWHM and energy resolution results of each full-energy peak are plotted in Fig.
 4.
-For Type-HEX06 modules, the energy resolution is 9% at 511 keV, which is noticeably better than the resolution achieved by the calorimeter used in the MACS experiment [41].
-Future improvements could include the adoption of large-area SiPM arrays with higher PDE.
+For the Type-HEX06 modules, the energy resolution is 10.8% at 511 keV, which is 6–7 times better than the resolution achieved by the calorimeter used in the MACS experiment.
+Figure 4: Calibration by Geant4 simulation of Type-HEX06 module.
+The red line is the FWHM curve fit by simulation data (blue dots with error bars), while the green line is the fit energy resolution.
+The errors are introduced by fitting the full-energy peak.
 II.3 Signal and Backgrounds
@@
 The event selection criteria can be summarized as follows:
-  • Threshold cut: $E_{thr}>0.2~{}\text{MeV}$;
-  • Energy difference cut: $\Delta E_{recon}<0.1~{}\text{MeV}$;
-  • Angle cut: $\theta_{recon}>160^{\circ}$.
-These criteria stand for trigger threshold, energy difference, and angle of two reconstructed hits, respectively.
+  • Trigger threshold: $E_{\text{Thr}}>2\sigma$;
+  • Energy difference cut: $\Delta E_{\text{Recon}}<3\sigma$;
+  • Angle cut: $\theta_{\text{Recon}}>170^{\circ}$.
+These criteria stand for trigger threshold, energy difference, and angle of two reconstructed tracks, respectively.
 Finally, the ECAL response spectra are smeared using the FWHM curve described by Eq.
@@
 Events that fall within the $3\sigma$ interval of the full-energy peak are identified as signal events.
-Figure 5: Spectra of original and reconstructed annihilation signals of M-to-$\bar{\text{M}}$ conversion by simulation.
+Figure 5: Spectra of single Type-PEN module (blue line), single Type-HEX module (green line) and reconstructed annihilation signals (red line) of M-to-$\bar{\text{M}}$ conversion by simulation.
 Fig.
 5 summarizes the simulated spectra of the positron annihilation signal.
-In the original $\gamma$ spectrum, the 511 keV full-energy peak accounts for only about 30% of the total counts.
-This is attributed to some of the $\gamma$-rays scattering to adjacent modules after depositing part of their energy, which forms a Compton continuum.
-Additionally, a peak at approximately 80 keV appears in the single $\gamma$ spectrum.
-This peak is attributed to characteristic X-rays emitted by lead atoms in the MCP.
-As a result, the signal efficiency is significantly reduced.
+In the original single-module $\gamma$ spectrum, the 511 keV full-energy peak accounts for only 47.2% of the total counts.
+This is attributed to some $\gamma$-rays scattering to adjacent modules after depositing part of their energy, which forms a Compton continuum.
+As a result, the signal efficiency will be significantly reduced.
 However, by clustering the energy responses from each event, the total energy of the scattered photons can be reconstructed.
@@
 The expected signal is thus represented by the $3\sigma$ interval of 1.022 MeV.
-Table 2: Contributions of signal Efficiency.
+Table 3: Summary of factors contributing to the signal efficiency for double $\gamma$ events of the ECAL.
 Items Efficiency
-Geometry 97%
-No cut 55.5%
-Energy difference cut 55.3%
-Threshold cut 49.9%
-Angle cut 43.5%
-Total 42.8%
-The result of the simulation indicates a 42.8%-55.5% signal efficiency in $3\sigma$ for ECAL.
-Losses of signal efficiency are as detailed in Table 2.
-Applications of various event selection criteria will depend on the real background level.
-The energy resolution and signal efficiency of ECAL are both superior to those of the MACS calorimeter, which is expected to significantly enhance the sensitivity of MACE.
-II.3.2 Beam Background
+Geometric efficiency of single $\gamma$ event $\varepsilon^{1\gamma}_{\text{Geom}}$ 95.2%
+Geometric efficiency of double $\gamma$ events $\varepsilon^{2\gamma}_{\text{Geom}}$ 93.6%
+Energy over threshold $\varepsilon_{\text{Trig}}$ ($E_{\text{Thr}}>2\sigma$) 98.7%
+Reconstruction efficiency $\varepsilon_{\text{Recon}}$ $\Delta E_{\text{Recon}}<3\sigma$ 90.8%
+$\theta_{\text{Recon}}>170^{\circ}$ 86.0%
+Total signal efficiency 78.3%
+The result of the simulation indicates a 78.3% signal efficiency for double $\gamma$ events in total.
+The factors contributing to the total signal efficiency are summarized in Table 3, where the standard deviation in the full-energy peak $\sigma$ is determined by the energy resolution.
+These results suggest that the efficiency performance of the ECAL surpasses that of the MACS calorimeter, which achieves an $\varepsilon^{1\gamma}$ of 79%.
+II.3.2 Beam-related backgrounds
 The muon beam, produced by stopping $\pi^{+}$ on the surface of a target, often includes positrons or other particles with similar momentum [54].
@@
 The beam positron background level is estimated based on this data.
-Assuming a $\mu^{+}$ beam flux of $10^{8}/$s, and a coincidence probability of $10^{-4}$ with the magnetic spectrometer, the upper limit of the background rates is established at 0.007/0.007/0.007 /year at 90% C.L.
-Thus, the ECAL is expected to be tolerant of both continuous and pulsed beams.
-II.3.3 Cosmic-Ray Muon Background
+Assuming a $\mu^{+}$ beam flux of $10^{8}/$s, and a coincidence probability of $10^{-4}$ with the magnetic spectrometer, the upper limit of the background rates is calculated to be $0.007/(10^{8}~{}\mu^{+}/\text{s}\cdot 1~{}\text{yr})$ at 90% C.L.
+II.3.3 Cosmic-Ray Muon backgrounds
 Cosmic-ray muons are another source of background in ECAL.
@@
 This includes some events where $e^{+}e^{-}$ annihilation produces 511 keV $\gamma$ rays.
-After applying coincidence selection, no events remain, resulting in background rates of less than 2/2/2 /year when considering coincidences with the Magnetic Spectrometer.
+After applying coincidence selection, no events remain, resulting in background rates of less than 2/2/2 /yr when considering coincidences with the Magnetic Spectrometer.
 Typically, a veto system can be installed outside ECAL to further reject cosmic-ray muon background [6, 57].
@@
 III.1 Experiment Setups and Results
-To characterize the parameters of the CsI(Tl) scintillator, a benchmark detector was constructed and calibrated using various radioactive sources at Sun Yat-sen University.
-It consists of a $3\times 3\times 8~{}\text{cm}^{3}$ CsI(Tl) crystal and a 1.5” PMT (HAMAMATSU CR284 with E2183-500 socket assembly).
+To characterize the parameters of the CsI(Tl) scintillator, a benchmark detector was constructed and calibrated using various radioactive sources at Sun Yat-sen University (SYSU).
+It consists of a $3\times 3\times 8~{}\text{cm}^{3}$ CsI(Tl) crystal and a 1.5-inch PMT (HAMAMATSU CR284 with E2183-500 socket assembly).
 The crystal is wrapped with ESR (glued) and coupled with PMT using optical grease, finally coated with 3M tape (Fig.
 6b).
-The readout system uses a Lecroy HDO4054A oscilloscope (Fig.
+The readout system uses a Lecroy HDO4054A oscilloscope (12 bit, 10 GS/s, see Fig.
 6c) for signal triggering, integration, and digitization.
-(a) CsI(Tl) crystal.
-(b) Benchmark detector.
-(c) Lecroy HDO4054A with a typical waveform of CsI(Tl) on screen.
-(d) CAEN V1730SB digitizer and the VMEbus.
-Figure 6: Photographs of the experiment setups.
-Figure 7: Measured spectra of ${}^{137}\text{Cs}$ (in green), ${}^{152}\text{Eu}$ (in red) at SYSU and ${}^{22}\text{Na}$ (in blue) at IMP.
-The benchmark detector is calibrated using ${}^{137}\text{Cs}$ and ${}^{152}\text{Eu}$ radioactive sources.
+The benchmark detector is calibrated using ${}^{137}\text{Cs}$ ($7.81\times 10^{3}$ Bq) and ${}^{152}\text{Eu}$ ($3.55\times 10^{4}$ Bq) radioactive sources.
 Radioactive sources were positioned close to the detector, and data acquisition was performed for 1800 s.
-When the PMT pulse signal exceeds the threshold, the oscilloscope triggers, integrates the waveform and records the charge value.
+When the PMT pulse signal exceeds the threshold, the oscilloscope triggers, integrates the waveform over a 2 $\mu$s time window and records the charge value.
 The calibration used the 662 keV peak from ${}^{137}\text{Cs}$ and the 40 keV, 122 keV, 244 keV, 344 keV, and 1408 keV peaks from ${}^{152}\text{Eu}$, as shown in Fig.
@@
 The spectra display the full-energy peaks of both isotopes.
-Using the FWHM curve, the energy resolution at 511 keV is calculated to be 11.5%.
-A more precise experiment was conducted to investigate the energy resolution at the Institute of Modern Physics, Chinese Academy of Sciences in Lanzhou, China.
+The energy resolution at 511 keV is calculated to be 11.5% by fitting the FWHM curve.
+(a) CsI(Tl) crystal.
+(b) Benchmark detector.
+(c) Lecroy HDO4054A displaying a typical waveform (pulse width of $\sim 2~{}\mu$s) of CsI(Tl).
+(d) CAEN V1730SB digitizer and the VMEbus.
+Figure 6: Photographs of the experiment setups.
+Figure 7: Measured spectra of ${}^{137}\text{Cs}$ (green line), ${}^{152}\text{Eu}$ (red line) at SYSU and ${}^{22}\text{Na}$ (blue line) at IMP.
+A more precise experiment was conducted to investigate the energy resolution at the Institute of Modern Physics (IMP), Chinese Academy of Sciences in Lanzhou, China.
 The absolute gain of the CR284 PMT was accurately calibrated in advance using single photoelectron (SPE) response measurements.
@@
 A 120 s measurement scans the high voltage ranging from 850 V to 1250 V.
-The PMT response waveforms are read out and recorded in ROOT files using the CAEN V1730SB digitizer (Fig.
+The PMT response waveforms are read out and recorded in ROOT files using the CAEN V1730SB digitizer (14 bit, 500 Ms/s, see Fig.
 6d), which is externally triggered by the waveform generator.
@@
 The CsI(Tl) crystal is wrapped with materials including Teflon, ESR (non-glued), and Tyvek.
-Other experiment setups were the same as Fig.
-6b, with ${}^{22}\text{Na}$ and ${}^{137}\text{Cs}$ sources used.
-Measurements were conducted for 600 s with the PMT operating at 850 V.
+Other experimental setups remained identical to Fig.
+6b, except for the use of ${}^{22}\text{Na}$ ($1.48\times 10^{5}$ Bq) and ${}^{137}\text{Cs}$ ($4.07\times 10^{4}$ Bq) sources.
+Measurements were conducted for 600 s with an integration time window of $2~{}\mu$s, under 850 V operating voltage of PMT.
 The energy resolution is directly calculated from the FWHM, and the results are listed in Table 4.
@@
 Moreover, the air gap between the non-glued reflectors and the crystal enhances the collection efficiency of scintillation photons through total internal reflection.
-In summary, Teflon is a promising reflector material for enhancing the energy resolution of MACE ECAL.
+(a) ${}^{22}\text{Na}$ source.
+(b) ${}^{137}\text{Cs}$ source.
+Figure 8: The energy spectra of ${}^{22}\text{Na}$ and ${}^{137}\text{Cs}$ in experiment (dots with error bars) and simulation (red line).
 This study primarily relies on Geant4 simulation for the design of the MACE calorimeter, requiring verification of the reliability of various physical processes in the simulation and a comparison with experimental results.
@@
 The CsI(Tl) crystal, Teflon reflector, optical grease, and CR284 PMT volume are modeled identically to the experimental setup.
-At the end of the simulation, the number of produced photoelectrons is recorded in a ROOT file event by event.
-The energy resolution is 8.2% at 511 keV in the simulation, while the experiment result is 10.3% (see Table 3).
-Figure 8: The photoelectron spectra of ${}^{137}\text{Cs}$ in experiment and simulation.
-The difference in energy resolution between the experiment and the simulation is reasonable.
 Fig.
-8 shows the spectra of photoelectrons reconstructed from experimental data and those recorded in simulation using ${}^{137}\text{Cs}$ source.
-Considering the mean photoelectron number of the 662 keV peak, the Geant4 simulation results are generally consistent with the experiment.
-However, there are still minor, predictable discrepancies.
-Several possible explanations are as follows.
-Firstly, the PMT gain value measurement may contain deviations caused by electronic noise or statistical errors.
-As a result, the number of photoelectrons reconstructed from the waveform is biased by approximately 76 PEs compared with the simulation.
-Secondly, the ideal simulation did not account for the natural radioactive background, resulting in a lack of response in the higher energy region.
-However, this contribution is expected to be limited due to the coincidence and reconstruction.
-The comparison between the experiment and the simulation of the benchmark detector gives the expected results with little divergence, which validates the design of the MACE ECAL and provides a foundation for further research.
+8 shows the spectra of calibrated experimental data and those recorded in simulation using ${}^{22}\text{Na}$ and ${}^{137}\text{Cs}$ source.
+The energy resolution is 9.7% at 511 keV in the simulation, compared to 10.3% in the experiment (see Table 4), resulting in a relative difference of 6%.
+Note that the difference in the width and height of the full-energy peak between the MC and the data originates from the discrepancy in energy resolution.
+Other minor discrepancies in the spectra shape are mainly likely caused by the insufficient modeling of electronic noise and nonlinearity, natural backgrounds, materials around the detector, and the simplified optical processes in Geant4 [59].
+This work focuses on a bench test of the CsI(Tl) crystal.
+However, important factors mentioned above that are difficult to simulate will be further studied in future prototype development efforts.
+In addition, we tested the SiPM readout using the NDL EQR15 11-6060D-S ($6\times 6~{}\text{mm}^{2}$) with a bias voltage set to 42 V in the Teflon reflector case.
+Using the same approach, an energy resolution of 26.1% is obtained.
+This significantly worse performance is attributed to the much smaller sensitive area of a single SiPM compared to the PMT.
+Simulation with the same configuration indicates that using a $4\times 4$ SiPM array with a total area of $24\times 24~{}\text{mm}^{2}$ could approximately double the number of collected photoelectrons, resulting in an energy resolution of 6.58% at 511 keV.
+The comparison between the experiment and the simulation of the benchmark detector gives the expected results of energy resolution at 511 keV, which validates the design of the MACE ECAL and provides a foundation for further research.
+The SiPM array readout, with its high PDE, low voltage requirements, and tolerance to magnetic fields, represents a promising upgrade option for the future.
 IV Conclusions and Perspectives
-After more than two decades of development of high-intensity muon source and detector technique, the first proposed muon-CLFV experiment in China, MACE, is expected to enhance the sensitivity of M-to-$\bar{\text{M}}$ conversion by three orders of magnitude.
+After more than two decades of development of high-intensity muon source and detector technique, the first proposed muon-CLFV experiment in China, MACE, is expected to enhance the sensitivity of M-to-$\bar{\text{M}}$ conversion by more than two orders of magnitude.
 This study focuses on the offline-software-based design of the $4\pi$ ECAL geometry for MACE.
 The geometry parameters were optimized using the Monte-Carlo simulation method, and the performance such as energy resolution, signal efficiency, and background level were evaluated.
-Simulation results indicate an energy resolution of 9% at 511 keV, and a signal efficiency of 55.5% for annihilation $\gamma$-ray events at maximum, significantly outperforming the MACS calorimeter design.
+Simulation results indicate an energy resolution of 10.8% at 511 keV, and a signal efficiency of 78.3% for annihilation $\gamma$-ray events, greatly outperforming the MACS calorimeter design.
 This research establishes a foundation for the subsequent technical development of the ECAL and supports the conceptual design of MACE.

### v4

@@
 The proposed MACE detector consists of a Michel electron Magnetic Spectrometer (MMS), a Positron Transport System (PTS), a Microchannel Plate (MCP) [24, 25], and an Electromagnetic Calorimeter (ECAL) to probe the conversion signal through final state coincidence.
-The specifications of muon beamline in MACE is expected to produce a surface muon flux on target of $10^{8}~{}\mu^{+}/$s, with a repetition rate of 30–50 kHz and a pulse width duty cycle below 20%.
+The specifications of muon beamline in MACE are expected to produce a surface muon flux on target of $10^{8}~{}\mu^{+}/$s, with a repetition rate of 30–50 kHz and a pulse width duty cycle below 20%.
 A schematic of the MACE apparatus is shown in Fig.
@@
 On the other hand, the atomic-shell positron will be accelerated by an electric field and guided through the PTS.
-The transverse position of the $e^{+}$ is conserved during the transportation to MCP, with a flight time FWHM of 6.9 ns.
-The cylindrical magnet with a field strength of 0.1 T outside ECAL is used to constrain the positron beam spot after transportation.
-While the MCP, with a timing resolution of 500 ps, is hit by a positron, the ECAL will be triggered and measure the produced two back-to-back annihilation $\gamma$-rays.
+The transverse position of the $e^{+}$ is conserved during the transportation to MCP, with a flight time full width at half maximum (FWHM) of 6.9 ns.
+The cylindrical magnet with a field strength of 0.1 T outside the ECAL is used to constrain the positron beam spot after transportation.
+When the MCP, with a timing resolution of 500 ps, is hit by a positron, the ECAL will be triggered and measure the produced two back-to-back annihilation $\gamma$-rays.
 As a rare-process searching experiment, MACE requires high sensitivity for the annihilation $\gamma$-rays to discriminate the signal from backgrounds and retain as much event information as possible.
@@
 Table 1 summarizes the properties of these inorganic scintillators.
-The selection of scintillation crystal for MACE ECAL is based on the energy region of interest less than 1 MeV.
+The scintillation crystal selection for the MACE ECAL is based on the energy region of interest less than 1 MeV.
 NaI(Tl) was widely used in the past decades due to its low cost and high scintillation light output.
@@
 Moreover, considering the overall budget of this experiment, ultrafast and bright crystals such as LYSO(Ce), GAGG(Ce), or LaBr3(Ce) are not currently included but may be considered for potential future upgrades.
-The low light output of PbWO4 leads to insufficient energy resolution in the low-energy regime.
+The low light output of PbWO4 leads to an insufficient energy resolution in the low-energy regime.
 The same is true for pure CsI.
-In the MACS experiment, the energy resolution of the pure CsI calorimeter for single $\gamma$ event and double $\gamma$ events was 69% and 52%, respectively [44, 14].
+In the MACS experiment, the energy resolution of the pure CsI calorimeter for single-$\gamma$ event and double-$\gamma$ events was 69% and 52%, respectively [44, 14].
 Doping thallium (Tl) in pure CsI can significantly increase its light yield, making it one of the brightest scintillators known.
-The objective of MACE ECAL is is to achieve excellent energy resolution for $\gamma$-rays produced in $e^{+}e^{-}$ annihilation.
+The objective of MACE ECAL is to achieve excellent energy resolution for $\gamma$-rays produced in $e^{+}e^{-}$ annihilation.
 This requires optimizing the light yield of the scintillator, which is a key factor in precise energy measurements.
@@
 The geometric parameters are ultimately chosen to be a crystal length of 10 cm and an ECAL inner radius of 20 cm based on consideration of both F.o.M and spatial constraints.
-In the context of the MACE experiment, the signal positrons from $\bar{\text{M}}$ decay are transported through a solenoid into the region of positron detection system.
+In the context of the MACE experiment, the signal positrons from $\bar{\text{M}}$ decay are transported through a solenoid into the region of the positron detection system.
 Some modules must be removed to ensure that the incoming signal positrons are not obstructed.
@@
 For an estimation of energy resolution, the class G4RadioactiveDecayPhysics is used to simulate the calibration of the Type-HEX06 module with radioactive sources, including 22Na, 137Cs, and 152Eu.
-Instead of the fully-covered SD in pervious simulation, the crystal is coupled with an 1.5-inch diameter PMT volume.
-An isotropic point source, wrapped in 1 mm thick plastic, is placed at the entrance of the module.
+Instead of the fully-covered SD in the previous simulation, the crystal is coupled with a 1.5-inch diameter PMT volume.
+An isotropic point source, wrapped in 1 mm thick plastic, is placed 5 cm away from the module.
 The energy dependence of FWHM is fit using Eq.
@@
 4.
-For the Type-HEX06 modules, the energy resolution is 10.8% at 511 keV, which is 6–7 times better than the resolution achieved by the calorimeter used in the MACS experiment.
+For the Type-HEX06 modules, the energy resolution is 10.8% at 511 keV, which is 6–7 times better than the resolution achieved by the calorimeter used in the MACS experiment [14].
 Figure 4: Calibration by Geant4 simulation of Type-HEX06 module.
@@
 Events that fall within the $3\sigma$ interval of the full-energy peak are identified as signal events.
-Figure 5: Spectra of single Type-PEN module (blue line), single Type-HEX module (green line) and reconstructed annihilation signals (red line) of M-to-$\bar{\text{M}}$ conversion by simulation.
+Figure 5: Spectra of a single Type-PEN module (blue line), a single Type-HEX module (green line) and reconstructed annihilation signals (red line) of M-to-$\bar{\text{M}}$ conversion by simulation.
 Fig.
@@
 The muon beam, produced by stopping $\pi^{+}$ on the surface of a target, often includes positrons or other particles with similar momentum [54].
-These background particles can enter the ECAL chamber in the same manner as the signal positrons, potentially causing accidental coincidences between MCP and ECAL.
+These background particles can enter the ECAL chamber in the same manner as the signal positrons, potentially causing accidental coincidences between the MCP and the ECAL.
 Therefore, it is crucial to estimate their potential contribution.
@@
 The EcoMug generator [56] is integrated into the MACE offline software to simulate cosmic-ray muon events on various types of surfaces (plane, cylinder, and half-sphere) with accurate angle and momentum distributions.
-To evaluate cosmic-ray muon background level in a potential future environment for the MACE detector, a hypothetical 18 m deep tunnel has been modeled in Geant4.
-A concrete volume, represented by G4Box with dimensions $50\times 50\times 18~{}\text{m}^{3}$, is created, with an air-filled cylindrical region, represented by G4Tubs, subtracted from its bottom to form a tunnel with a 1 m radius and 10 m length.
+To evaluate the cosmic-ray muon background level in a potential future environment for the MACE detector, a hypothetical 18 m deep tunnel has been modeled in Geant4.
+A concrete volume, represented by G4Box with dimensions $50\times 50\times 18~{}\text{m}^{3}$, is created, with an air-filled cylindrical region, represented by G4Tubs, subtracted from its bottom to form a tunnel with a 1 m radius and a 10 m length.
 ECAL is positioned at the center of the tunnel, surrounded by a copper solenoid and an iron shield.
@@
 This includes some events where $e^{+}e^{-}$ annihilation produces 511 keV $\gamma$ rays.
-After applying coincidence selection, no events remain, resulting in background rates of less than 2/2/2 /yr when considering coincidences with the Magnetic Spectrometer.
+After applying coincidence selection, no events remain, resulting in a background rate of less than 2/2/2 /yr when considering coincidences with the Magnetic Spectrometer.
 Typically, a veto system can be installed outside ECAL to further reject cosmic-ray muon background [6, 57].
+II.3.4 Radioactive contamination background
+It is worth mentioning that studies have reported the radioactive contamination of scintillation materials (e.g., 137Cs in CsI), which may greatly affect low-counting experiments.
+In the context of the MACE PDS, if the MCP is triggered by its dark counts while the radionuclides within the ECAL materials decay simultaneously, a false coincidence may occur.
+Based on the provided data, a single CsI(Tl) module in the ECAL will have a 662 keV $\gamma$-ray background counting rate up to 0.07 Hz [58].
+A preliminary estimation of the accidental coincidence rate can be conducted by
+$\begin{gathered}n=2\tau n_{\text{MCP}}n_{\text{ECAL}}~{},\\ n_{\text{ECAL}}=2\tau n_{\text{mod1}}n_{\text{mod2}}\times N_{\text{mod}}/2~{},\end{gathered}$ (3)
+where $\tau$ is the coincidence resolving time which is set to be CsI(Tl) decay time, $n_{\text{MCP}}$ is the dark count rate of the MCP (commonly $\sim 1$ kHz), $n_{\text{ECAL}}$ is the accidental coincidence of two modules caused by the natural background detected by the ECAL and $N_{\text{mod}}$ is the total module numbers.
+Eventually, the rate is calculated to be 0.19/0.19/0.19 /yr.
+This background, however, falls outside the $3\sigma$ interval of the signal of interest.
+If the event selection criteria are applied, the background caused by dark count and natural radioactivity can be neglected.
 III Validation with a Benchmark Detector
@@
 The benchmark detector is calibrated using ${}^{137}\text{Cs}$ ($7.81\times 10^{3}$ Bq) and ${}^{152}\text{Eu}$ ($3.55\times 10^{4}$ Bq) radioactive sources.
-Radioactive sources were positioned close to the detector, and data acquisition was performed for 1800 s.
-When the PMT pulse signal exceeds the threshold, the oscilloscope triggers, integrates the waveform over a 2 $\mu$s time window and records the charge value.
+Radioactive sources were positioned 5 cm away from the detector, and data acquisition was performed for 1800 s.
+When the PMT pulse signal exceeds the threshold, the oscilloscope triggers, integrates the waveform over a 2 $\mu$s time window, and records the charge value.
 The calibration used the 662 keV peak from ${}^{137}\text{Cs}$ and the 40 keV, 122 keV, 244 keV, 344 keV, and 1408 keV peaks from ${}^{152}\text{Eu}$, as shown in Fig.
@@
 (b) ${}^{137}\text{Cs}$ source.
-Figure 8: The energy spectra of ${}^{22}\text{Na}$ and ${}^{137}\text{Cs}$ in experiment (dots with error bars) and simulation (red line).
+Figure 8: The energy spectra of ${}^{22}\text{Na}$ and ${}^{137}\text{Cs}$ in the experiment (dots with error bars) and the simulation (red line).
 This study primarily relies on Geant4 simulation for the design of the MACE calorimeter, requiring verification of the reliability of various physical processes in the simulation and a comparison with experimental results.
@@
 The CsI(Tl) crystal, Teflon reflector, optical grease, and CR284 PMT volume are modeled identically to the experimental setup.
+The simulation takes into account the natural background contributions from 238U and 40K.
 Fig.
@@
 This work focuses on a bench test of the CsI(Tl) crystal.
-However, important factors mentioned above that are difficult to simulate will be further studied in future prototype development efforts.
+Therefore, important factors mentioned above that are difficult to simulate will be further studied in future prototype development efforts.
 In addition, we tested the SiPM readout using the NDL EQR15 11-6060D-S ($6\times 6~{}\text{mm}^{2}$) with a bias voltage set to 42 V in the Teflon reflector case.
@@
 This research establishes a foundation for the subsequent technical development of the ECAL and supports the conceptual design of MACE.
-However, the studies on accidental coincidence backgrounds from the beam and cosmic rays are currently idealized, and the effects of electronic noise and natural radioactive backgrounds have not yet been evaluated.
 A prototype of ECAL will be constructed for additional validation in progress in these areas.